Exercise 08-3 Preparing flexible budgets LO P1 Tempo Companys fixed budget (based on sales of 14,000 units) for the first quExercise 08-3 Preparing flexible budgets LO P1 Tempo Companys fixed budget (based on sales of 14,000 units) for the first quExercise 08-3 Preparing flexible budgets LO P1 Tempo Companys fixed budget (based on sales of 14,000 units) for the first quExercise 08-3 Preparing flexible budgets LO P1 Tempo Companys fixed budget (based on sales of 14,000 units) for the first qu
Answer #1
Variable Amount
per Unit
Total Fixed Cost
Sales 209.00 =2926000/14000
Variable costs:
Direct materials 25.00 =350000/14000
Direct labor 43.00 =602000/14000
Production supplies 27.00 =378000/14000
Sales commissions 9.00 =126000/14000
Packaging 15.00 =210000/14000
Total variable costs 119.00
Contribution margin 90.00
Fixed costs:
Plant manager salary 150,000
Advertising 100,000
Administrative salaries 200,000
Depreciation—office equip. 170,000
Insurance 140,000
Office rent 150,000
Total fixed costs 910,000
1
Variable cost per unit 119
2
Total fixed costs 910000
3
Income from Operations at sales of 12000 units 170000 =(12000*90)-910000
4
Income from Operations at sales of 16000 units 530000 =(16000*90)-910000